# Neuromodulation and homeostasis: complementary mechanisms for robust neural function

### v1

@@ DISCUSSION
$[g_{\mathrm{f}}(V);g_{\mathrm{s}}(V);g_{\mathrm{u}}(V)]=f_{\mathrm{DIC}}(V)=S(V)\cdot\bar{g}_{\mathrm{ion}}\,,$
where $S(V)\in\mathbb{R}^{3\times N}$ is a sensitivity matrix that can be built by: $S_{i\,j}(V)=-\left(w_{i\,j}\cdot\frac{\partial\dot{V}}{\partial X_{j}}\frac{\partial X_{j,\infty}}{\partial V}\right)/g_{\mathrm{leak}}$, where $i$ denotes the timescale, $X_{j}$ are gating variables of the j-th channel of the considered model and $w_{i\,j}$ is a timescale-dependent weight which is computed as the logarithmic distance of the time constant of $X_{j}$ and the timescale $i$ (Drion et al., 2015a, ). While the complete curve of the DICs may be of interest, only its value at the threshold voltage $V_{\mathrm{th}}$ is used, as the values and signs of the DICs at $V_{\mathrm{th}}$ reliably determine the firing pattern Drion et al., 2015a . Thus, the following linear system $f_{\mathrm{DIC}}(V_{\mathrm{th}})=S(V_{\mathrm{th}})\cdot\bar{g}_{\mathrm{ion}}$ makes the link between ion channel conductances and neuronal activity.
COMPETING INTEREST
The authors declare no competing interests.
FUNDING
Arthur Fyon is a Research Fellow of the “Fonds de la Recherche Scientifique - FNRS”, supported by grant ASP-REN40024838. This work was supported by the Belgian Government through the Federal Public Service (FPS) Policy and Support (BOSA) grant NEMODEI (NEuroMOrphic Design of Embodied Intelligence).
AUTHOR CONTRIBUTIONS STATEMENT
A.F. and G.D. conceived and designed the research. A.F. and G.D. developed the model code. A.F. performed the simulations and analyzed the data. A.F. and G.D. interpreted the results. A.F. prepared figures and drafted the manuscript. A.F. and G.D. edited and revised the manuscript. A.F. and G.D. approved the final version of the manuscript.
DATA AVAILABILITY
All code and data can be found on the first author’s GitHub (https://github.com/arthur-fyon/BIOCONTROL_2025).

### v2

@@ Discussion
$g_{\mathrm{f}}(V);g_{\mathrm{s}}(V);g_{\mathrm{u}}(V)]=f_{\mathrm{DIC}}(V)=S(V)\cdot\bar{g}_{\mathrm{ion}}\,$,
where $S(V)\in\mathbb{R}^{3\times N}$ is a sensitivity matrix that can be built by: $S_{i\,j}(V)=-\left(w_{i\,j}\cdot\frac{\partial\dot{V}}{\partial X_{j}}\frac{\partial X_{j,\infty}}{\partial V}\right)/g_{\mathrm{leak}}$, where $i$ denotes the timescale, $X_{j}$ are gating variables of the j-th channel of the considered model and $w_{i\,j}$ is a timescale-dependent weight which is computed as the logarithmic distance of the time constant of $X_{j}$ and the timescale $i$ [10]. While the complete curve of the DICs may be of interest, only its value at the threshold voltage $V_{\mathrm{th}}$ is used, as the values and signs of the DICs at $V_{\mathrm{th}}$ reliably determine the firing pattern [10, 4]. Thus, the following linear system $f_{\mathrm{DIC}}(V_{\mathrm{th}})=S(V_{\mathrm{th}})\cdot\bar{g}_{\mathrm{ion}}$ makes the link between ion channel conductances and neuronal activity.
Use of generative AI
The ChatGPT-5 chatbot has been used to improve the syntax and grammar of several paragraphs in the manuscript. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Supporting information
S1 Appendix
Details and additional experiments We provide all the elements required to reproduce the results presented in this paper, as well as additional experiments and their results. This includes, for example, model equations and their parameter values.